# The Fall of ROME: Understanding the Collapse of LLMs in Model Editing

### v1

@@ 4 A Simple Solution to Avoid Collapse
Having identified the reasons for ROME’s collapse, it is crucial to provide a solution to prevent these problems. C-ROME introduced in § 3.1 can effectively keep the stability of edited models, but Table 3 reveals that it fails to successfully integrate target knowledge into the model, as evidenced by its low efficacy and generalization Yao et al. (2023) metrics on collapse cases. The reason is that C-ROME employs prefixed keys $k_{*}$ only when editing. However, during the evaluation of collapse cases, the prompts used to assess efficacy adopt unprefixed keys $k_{*}^{s}$, which significantly differs from $k_{*}$. This inconsistency results in an inability to obtain the appropriate target value vector corresponding to the key of collapse cases, finally leading to a failure in efficacy.
To address this issue, we propose a straightforward solution, which appends a random prefix during the testing phase to the prompt for cases where the key corresponds to the first token. The results in Table 4 demonstrate that this method significantly raises the efficacy for both GPT-2-XL and GPT-J, albeit with a relatively limited improvement of generalization. The suboptimal performance on the collapse cases of Llama2-7b is due to their different pattern from that observed in other two models. Nonetheless, such cases are extremely rare ($21$ out of $21919$ in the COUNTERFACT dataset), and their collapse has effectively been avoided.
Model Cases efficacy generalization locality GPT-2-XL collapse $100$% $16.88$% $100$% normal $96.16$% $41.88$% $97.34$% GPT-J collapse $100$% $32.94$% $89.41$% normal $99.77$% $50.00$% $95.61$% Llama2-7b collapse $12.70$% $12.70$% $100$% normal $91.95$% $46.73$% $97.56$%
@@ 5 Conclusion and Future Work
In this paper, we conduct a thorough investigation into the underlying causes of LLM’s collapse triggered by a single edit of ROME. Our extensive experiments demonstrate that such collapse arises from two aspects: i) irregularities in the official implementation of ROME, which employs two types of keys in parameter updating; ii) anomalous distribution of the first token in GPT-like models. Consequently, we propose a straightforward method to address the model collapse issue of ROME, and conduct experiments to validate its effectiveness.
@@ A.1 Distribution of Keys in Other LLMs
The distribution of $C^{-1}k_{*}$ and $k_{*}^{s}$ for collapse and normal cases of GPT-J in two-dimensional space is shown in Figure 4(a), demonstrating a significant difference between the distributions of these two elements in collapse cases. The results for $k_{*}$ and $k_{*}^{s}$ is depicted in Figure 4(b), revealing similar disparities. The corresponding results for Llama2-7b are provided in Figure 5(a) and Figure 5(b), showing consistent phenomena.
A.2 First token in T5-3B
To explore whether the anomalous distribution of the first tokens in GPT-like models can be attributed to their inability to interact with subsequent tokens within autoregressive models, we take the encoder-decoder model T5-3B as an counterexample and observe the distribution of an equal number of first tokens compared to subsequent tokens across various layers in its encoder. The results in Figure 6 indicate that there is no significant difference between the representations of the first token and subsequent tokens, corroborating our hypothesis.
A.3 Llama2-7b without Prepended Token
We manually removed the prepended token <s> in Llama2-7b, thereby positioning the key $k_{*}^{s}$ of the collapse case as the first token of the input. In this setting, we employed ROME to edit Llama2-7b on the collapse cases of GPT-2-XL. The results presented in Figure 3 indicate that Llama2-7b also succumbs to collapse after editing.
Figure 3: Scatter plot of perplexity for Llama2-7b models edited by ROME, with each point representing a unique edit case in the collapse case of GPT-2-XL. “Case ID” refers to the index of each edit sample.

### v2

@@ 4 A Simple Solution to Avoid Collapse
Having identified the reasons for ROME’s collapse, it is crucial to provide a solution to prevent these problems. C-ROME introduced in § 3.1 can effectively keep the stability of edited models, but Table 4 reveals that it fails to successfully integrate target knowledge into the model, as evidenced by its low efficacy and generalization Yao et al. (2023) metrics on collapse cases. This failure arises from the inconsistency of C-ROME between editing and testing. Specifically, C-ROME employs prefixed keys $\overline{\bm{k}}$ only when editing, while during testing, the prompts used to evaluate efficacy adopt unprefixed keys $\bm{k}^{u}$, which significantly differ from $\overline{\bm{k}}$. This inconsistency results in an inability to obtain the appropriate target value vector corresponding to the key of collapse cases, finally leading to low efficacy of editing.
To address this issue, we propose a straightforward solution, which appends a random prefix, drawn from those utilized in the editing process, to the prompt of collapse cases during the testing phase. The results in Table 5 demonstrate that this method significantly improves the efficacy for GPT-2-XL, GPT-J, and Llama2-7b, albeit with a relatively limited improvement of generalization.
Model Cases efficacy generalization locality GPT-2-XL collapse $100$% $16.88$% $100$% normal $96.16$% $41.88$% $97.34$% GPT-J collapse $100$% $32.94$% $89.41$% normal $99.77$% $50.00$% $95.61$% Llama2-7b collapse $91.27$% $29.37$% $100$% normal $91.95$% $46.73$% $97.56$%
@@ 5 Conclusion and Future Work
In this paper, we thoroughly investigate the underlying causes of LLMs collapse triggered by a single edit of ROME. Our extensive experiments demonstrate that such collapse arises from two aspects: i) irregularities in the official implementation of ROME, which employs two types of keys in parameter updating; ii) anomalous representation distribution of the first token in autoregressive models. Consequently, we propose a straightforward and simple method to address the model collapse issue of ROME, and validate its effectiveness with extensive experiments
@@
Acknowledgements
This work was supported by the National Key R&D Program of China (2022YFB3103700, 2022YFB3103704), the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0680101), and the Innovation Funding of ICT, CAS (E361120).
@@ A.1 Distribution of Keys in Other LLMs
The distribution of $C^{-1}\overline{\bm{k}}$ and $\bm{k}^{u}$ for collapse and normal cases of GPT-J in two-dimensional space is shown in Figure 4(a), demonstrating a significant difference between the distributions of these two elements in collapse cases. The results for $\overline{\bm{k}}$ and $\bm{k}^{u}$ is depicted in Figure 4(b), revealing similar disparities. The corresponding results for Llama2-7b are provided in Figure 5(a) and Figure 5(b), showing consistent phenomena.
A.2 Results without Prepended Token
@@
To validate that the absence of collapse in Llama2-7b, Mistral-7b, and Llama3-8b for the collapse cases of GPT-2-XL and GPT-J, is due to the addition of a prefix token, we manually removed the prepended token of these models, thereby positioning the unprefixed key $\bm{k}^{u}$ of the collapse cases as the first token of the input. In this setting, we employed ROME to edit these three models on the collapse cases of GPT-2-XL and GPT-J. The results presented in Figure 7 indicate that Llama2-7b, Mistral-7b, and Llama3-8b also succumb to collapse after editing.
A.3 Representation of First Token in T5-3B
Figure 7: Scatter plot of perplexity for Llama2-7b, Mistral-7b, and Llama3-8b models edited by ROME, with each point representing a unique edit case in the collapse cases of GPT-2-XL and GPT-J. “Case ID” refers to the index of each edit sample.
The anomalous representation distribution of the first tokens in autoregressive models may be attributed to their inability to interact with subsequent tokens. To verify it, we take the encoder-decoder model T5-3B as a counterexample and analyze the representation distribution of the first tokens in the collapse cases compared to an equal number ($77$) of subsequent tokens from the normal cases across various layers in its encoder. The results in Figure 6 indicate that there is no significant difference between the representations of the first token and subsequent tokens, corroborating our hypothesis.
A.4 Impact of Position Embedding
In this section, we conducted experiments on GPT-2-XL, GPT-J, and Llama2-7b to investigate whether the anomalous distribution of the first token is attributable to its position embedding. For Llama2-7b, we removed the special token <s> that the tokenizer additionally prepends at the beginning of the input to maintain consistency with GPT-2-XL and GPT-J.
Model Perplexity Original Second2First GPT-2-XL min $2177.82$ $1008.21$ avg $19877.79$ $1397.87$ max $179185.99$ $2153.86$ GPT-J min $5094.73$ $8153.70$ avg $28835.21$ $26978.14$ max $85936.24$ $124982.41$ Llama2-7b min $16279.75$ $17561.97$ avg $67436.51$ $72692.50$ max $206307.60$ $349577.58$
Table 6: The minimum, average, and maximum perplexity observed in collapse cases when utilizing the original position embeddings (Original) and when assigning the first token’s position embedding as that of the second token (Second2First) for various LLMs.
For collapse cases where the subjects are the first tokens, we set the position embedding of the first token as that of the second token (Noted as Second2First). The results presented in Table 6 indicate that this approach mitigates model collapse on GPT-2-XL, but it is completely ineffective on GPT-J and Llama2-7b.
Model Perplexity Original First2Second GPT-2-XL min $68.55$ $81.39$ avg $68.81$ $39714.90$ max $69.03$ $912001.20$ GPT-J min $48.80$ $48.47$ avg $49.03$ $48.68$ max $49.50$ $49.48$ Llama2-7b min $32.83$ $33.14$ avg $33.32$ $2104.90$ max $37.03$ $42154.10$
Table 7: The minimum, average, and maximum perplexity observed in normal cases when utilizing the original position embeddings (Original) and when assigning the second token’s position embedding as that of the first token (First2Second) for various LLMs.
For normal cases where the subjects are the second tokens, we assign the position embedding of the second token as that of the first token (Noted as First2Second). The results in Table 7 reveal that this change leads to partial model collapse in GPT-2-XL and Llama2-7b, but all edited models of GPT-J remain stable.
The results from the two aforementioned aspects suggest that position embedding may be a contributing factor to the abnormal representation of the first token, but it is not the sole factor.
A.5 Collapse of First Token Representation
From Figure 2 and Figure 3, we observed an unusual phenomenon that the collapse keys $\bm{k}^{u}$ (i.e., representations of the first tokens) appear to be more concentrated than the normal keys $\bm{k}^{u}$ (i.e., representations of the subsequent tokens). To assess the degree of aggregation of the first tokens and subsequent tokens, we calculated the average distance of each element from the cluster center for both the first tokens and all the subsequent tokens, denoted as $D\left(F\right)$ and $D\left(S\right)$, correspondingly.
$D=\frac{1}{N}\sum_{i=1}^{N}\left\|\bm{e}_{i}-\frac{1}{N}\sum_{k=1}^{N}\bm{e}_{k}\right\|_{2}$ (4)
Here, $\bm{e}_{i}$ and $\bm{e}_{k}$ represent the embeddings of the $i$-th and $k$-th tokens, which are the outputs of the first MLP layer within the transformer block.
With this metric established, we computed the values within the edited layers of GPT-2-XL, yielding $D\left(F\right)$ being $0.578$ and $D\left(S\right)$ being $13.895$. The result suggests a markedly higher concentration in the representations of the first tokens compared to those of subsequent tokens. This observation raises a further question: Given that different first tokens have distinct embeddings when input into the transformer, why are their representations in the middle layers so closely concentrated?
To investigate this, we computed the distances $D\left(F\right)$ and $D\left(S\right)$ from the first layer to the edited layer (layer 17) in GPT-2-XL. As depicted in Figure 8(a), prior to layer 8, $D\left(F\right)$ and $D\left(S\right)$ exhibit no significant divergence. However, post layer 8, the representations of the first tokens rapidly shrink. The same phenomenon is also observed in GPT-J, as shown in Figure 8(b). However, our experimental results indicate that such phenomenon does not appear on Llama2-7b, Mistral-7b, and Llama3-8b. Consequently, we decide not to delve further into this particular aspect.
The underlying causes of the first token’s representation concentration in GPT-2-XL and GPT-J remain unclear. A potential factor, as explored in Appendix A.3, is that within autoregressive LLMs, the first token cannot interact with subsequent tokens. Continuous self-interaction may lead to the contraction of its representation. Since this phenomenon is not related to the model collapse during editing examined in this paper, it has been remained for future research.
(a) GPT-2-XL
(b) GPT-J
Figure 8: Average distances of each element from the cluster center for the first tokens and the subsequent tokens, across layers from the first layer to the edited layer in GPT-2-XL and GPT-J.